Required information The Foundational 15 (LO2-1, LO2-2, LO2-3, LO2-4) [The following information applies to the questions dis6. What was the total manufacturing cost assigned to Job Q? (Do not round intermediate calculations.) Total manufacturing cos7. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answ8. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establi9. What was Sweeten Companys cost of goods sold for March? (Do not round intermediate calculations.) Cost of goods sold
Answer #1

Calculate Predetermined overhead Rate

Estimated total fixed manufacturing overhead $27,000
EstimatedVariable overhead cost of molding dept.

$4750

2500 x 1.90

Estimated Variable overhead cost of fabrication dept.

$4050

1500 x 2.70

Estimated Total Overhead $35800

Predetermined Overhead Rate =   Estimated total Overhead
Estimated Total machine hour

= 35800/4000

=8.95 per MH

Manufacturing overhead Applied (Job Q)

= Predetermined Overhead Rate *  Actual Machine hour

= 8.95 * 2700

= $ 24165

  

6.Total manufacturing cost assigned to job Q

Direct material $10500
Direct Labor $9500
Manufacturing Overhead $24165
Total manufacturing Cost 44165

7.Unit product cost of Job Q If it included 30 units

Total Manufacturing Cost (a) 44165
Total Units ( b) 30
Unit Cost a/b 1472

8..Total Selling Price & Price per Unit

Job P Job Q
Total Manufacturing Cost

72535

44165
Mark Up (80% of Total manufacturing Cost) 58028 35332
Total Price for the Job 130563 79497

Job p Total manufacturing Cost = 18000+ 25000+(8.95 * 3300) = 72535

Add 80% mark up to manufacturing cost calculated

Job P Job Q
Total price for the job (a) 130563 79497
Number of Units (b) 20 30
Unit Price a/b 6528 2650

9.Cost of Goods Sold in March

Total Manufacturing Cost assigned to Job P (a) 72535
Total Manufacturing Cost assigned to Job Q (b) 44165
Cost of Goods sold a + b 116700